OpenStudy (anonymous):

if sinx,cosx,tanx are in g.p then (cosx)^9+(cosx)^6+3(cosx)^5-1 equals

OpenStudy (mathmale):

Please clarify what you mean by "g. p."

OpenStudy (anonymous):

g.p. is geometric progression

OpenStudy (mathmale):

New situation for me. What does it signify when you state that sin x, cos x and tan x are in geometric progression?

OpenStudy (anonymous):

that is cosx=sinx*r and tanx=cosx*r where r is the common ratio

OpenStudy (anonymous):

also tanx=sinx*r^2 and \[\cos x =\sqrt{sinx*tanx}\]

OpenStudy (mathmale):

If you state that that is cosx=sinx*r and tanx=cosx*r where r is the common ratio, then it appears that the common ratio, r, is cos x / sin x, or cot x. Is that your reasoning?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

if that is the case, is it true, as you say, that tanx=sinx*r^2? Try verifying that.